# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v2

@@ Abstract
The “Last Mile Challenge” has long been considered an important, yet unsolved, challenge for autonomous vehicles, public service robots, and delivery robots. A central issue in this challenge is the ability of robots to navigate constrained and cluttered environments that have high agency (e.g., doorways, hallways, corridor intersections), often while competing for space with other robots and humans. We refer to these environments as “Social Mini-Games” (SMGs). Traditional navigation approaches designed for MRN do not perform well in SMGs, which has led to focused research on dedicated SMG solvers. However, publications on SMG navigation research make different assumptions (on centralized versus decentralized, observability, communication, cooperation, etc.), and have different objective functions (safety versus liveness). These assumptions and objectives are sometimes implicitly assumed or described informally. This makes it difficult to establish appropriate baselines for comparison in research papers, as well as making it difficult for practitioners to find the papers relevant to their concrete application. Such ad-hoc representation of the field also presents a barrier to new researchers wanting to start research in this area. SMG navigation research requires its own taxonomy, definitions, and evaluation protocols to guide effective research moving forward. This survey is the first to catalog SMG solvers using a well-defined and unified taxonomy and to classify existing methods accordingly. It also discusses the essential properties of SMG solvers, defines what SMGs are and how they appear in practice, outlines how to evaluate SMG solvers, and highlights the differences between SMG solvers and general navigation systems. The survey concludes with an overview of future directions and open challenges in the field.
Index Terms:
Multi-Agent Interactions, Dynamic Games, Social Navigation
I Introduction
@@
(b) General MRN (taken with permission from [23]).
@@
In recent years, the research field of MRN has started to coincide with the momentum of robot deployments in shared human spaces, largely driven by advances in data, compute, and foundation models. As a consequence, research has flourished in MRN, and more specifically, MRN in constrained environments that arise in shared human spaces. In Figure 1, we highlight two types of MRN in constrained scenarios. On the left, we depict Social Mini-Games (SMGs) [22, 24] on the right we depict “general MRN”. A key distinguishing factor between these scenarios is the notion of agency–that is, the capacity of an agent to take independent actions and influence the environment around them. SMGs exhibit high agency where small actions can significantly change the interaction dynamics, coordination requirements, and potential for deadlocks. General MRN, on the other hand, exhibit low to medium agency.
High-agency SMGs have several unique challenges. First, without explicit scheduling, robots result in deadlocks, collisions, or non-smooth trajectories [25, 26]. Second, humans can avoid collisions and deadlocks without having to deviate too much from their nominal speed or trajectory. For instance, when two individuals go through a doorway together, one person modulates their velocity by just enough to enable the other person to pass through first, while still adhering closely to their preferred speed. This type of behavior presents a significant challenge for robots, especially in high-agency SMGs.
Classical navigation approaches typically fail to cover all the aspects needed for SMGs [22]. So, new techniques, ideas, and systems are being investigated, often with roots deep within the larger parent field of MRN. This fact presents the crux of the issue: researchers studying SMG solvers from a MAPF perspective will build off of MAPF research, MARL-based SMG solvers will form their roots from the MARL community, and so on. Other communities build upon game theory and optimization. Each field has its own set of assumptions, limitations, objectives, and techniques. As a result, MRN in SMG research has fractured into these sub-communities without a clear, unified formulation. This categorization into sub-communities presents an entry barrier to new researchers in this field by making it difficult to navigate through and understand existing literature and to establish appropriate baselines for comparison. In addition, this makes it difficult for practitioners to find papers and solutions relevant to their concrete application.
This article aims to address this growing challenge by introducing a unified taxonomy to describe MRN in SMG problems, and by establishing common assumptions, properties, and measures for evaluating MRN in SMG algorithms. The unified taxonomy we present in this paper is our attempt to classify the currently studied SMG solver variants. We hope this terminology will serve as a common ground for future researchers, and will be used by them to describe their contributions succinctly and accurately.
I-A Differentiation from Existing Surveys
@@
Our survey focuses on multi-robot navigation in social mini-games, and thus belongs to the broader field of social robot navigation. Social robot navigation, however, is vast and we will not attempt to summarize it; instead, we refer to many recent surveys on social navigation [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Among these, [32] focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations on how to properly evaluate social navigation research. Another recent survey by Mavrogiannis et al. [33] focuses on the core challenges of social navigation with respect to different algorithms, human behavior models, and evaluation, but they focus on the general navigation case. Wang et al. [36] proposes new metrics evaluating the principles defined in [27], such as comfort, naturalness, and sociability. And finally, datasets [37, 38] and benchmarks [39, 40] have enabled researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. However, all of the surveys discussed above on various aspects of social robot navigation focus on robots in sparse or dense crowds, but not social mini-games. We showcase the difference in Figure 1.
Perhaps the closest survey that mirrors the same objective as with this survey (that is, attempting to unify a growing, but fractured, research area) is the seminal survey on MAPF (“Multi-Agent Pathfinding: Definitions, Variants, and Benchmarks”) by Stern et al. [41]. Here, the authors present a definition for MAPF, describe the different variants of it, and introduce two benchmarks for proper evaluation. This paper was influential in anchoring and streamlining the future research in MAPF. We envision a similar success via our survey.
The overall structure of this survey is given in Fig. 2. We first define the SMG problem in Sec. II, highlight its various scenarios where it appears, and how to evaluate navigation approaches in SMGs. Then in Sec. III, we introduce the taxonomy, properties, and algorithmic categories for research on SMG solvers. Then in Sec. IV, we benchmark and compare some recent state of the art SMG solvers on the scenarios introduced earlier. Finally in Sec. V, we conclude by discussing general trends, open problems, and future directions.
@@
II Social Mini-Games
II-A SMGs versus General MRN
@@
Below, we give a mathematical definition of SMGs. It basically says that if we assume each agent has one or more individual optimal trajectory (for example, in the doorway scenarios, the optimal trajectory for both the green and the orange agent would be a straight line path towards the red cross.), then if all the individual optimal trajectories intersect for all agents within some time interval, then that scenario is an SMG.
@@
SMGs are embedded within MRN: most trajectories are decoupled, but at specific times (bottlenecks, crossings, occlusions) the interaction becomes a game with coupled best responses. MAPF also resolves conflicts but (classically) in discrete, fully observed, centralized settings that compute joint plans offline; SMGs arise online, continuous-time, often decentralized, with heterogeneous objectives and social norms. Furthermore, it is typical for SMGs to include around 2-4 agents in the most commonly occurring SMGs. Of course, an environment may consist of multiple SMGs occurring simultaneously. Our scenarios (Sec. II-B) are chosen because they activate specific SMG characteristics, and our metrics (Sec. II-C) include SMG-aware outcomes that are immaterial in decoupled MRN segments but decisive inside SMGs.
@@ II-C Evaluation Metrics
III Algorithms, Taxonomy and Properties for SMGs
We analyze algorithms and system design choices conditioned on an active SMG, i.e., when at least one SMG property from Sec. II-B is present (bottleneck contention, crossing flows, persistent mutual occupancy, occlusion, or conflicting objectives). Let $\mathcal{K}(t)\!\subseteq\!\{1,\dots,N\}$ denote the active coupling set, that is, the minimal subset of agents whose best responses are mutually coupled during $[a,b]$ (Sec. II-B). Coordination, communication, observability, invasiveness, cooperation, and deadlock handling are interpreted on $\mathcal{K}(t)$, not on the full set of agents. Throughout, we tie design choices to game-aware outcomes introduced in Sec. II-C: Social Welfare Gap (SWG), Turn-Taking Fairness (TTF), and Invasiveness Score (IS), reported alongside standard MRN metrics. This section thus specifies how MRN methods become SMG solvers once a trigger activates and a coupling set emerges.
III-A Categorization of SMG Solvers by Paradigm
@@
III-A1 Multi-Agent Reinforcement Learning (MARL)
@@
III-A2 Multi-Agent Path Finding (MAPF)
@@ III-A4 Others
III-B Taxonomy of SMG Design Choices
@@ III-B1 Coordination
The last type of coordination is Decentralized which allows each agent to make its own decisions independently. Instead of relying on a central authority, each agent acts based on its own observations, rewards, and goals. This approach promotes independence and flexibility, enabling agents to adapt to their surroundings and other agents dynamically [77, 78]. Although it may seem less organized, with the right algorithms and planners, decentralized coordination can generate effective and efficient controls to develop a safe, collision-free, and deadlock-free trajectory [79, 44, 80, 81, 82]. Examples of this approach include IMPC-DR [6] and ORCA-MAPF [8], where agents independently navigate in their environment while avoiding collisions and deadlocks. Figures 5 and 6 list the key papers for the Centralized and Decentralized communication categories. Note that the color legend (from Fig. 4) signifies the categorization based on Algorithmic/Methodological paradigm discussed in Sec. III-A, where we use different colors for different approaches- multi-robot Reinforcement Learning (cyan), Optimization (violet), multi-robot Path Finding (green), and Others (Orange).
[…]
Figure 4: Color legend across figures: Categories of MRN Algorithms based on Algorithmic/Methodological Paradigm.
@@
III-B2 Communication
@@
III-B3 Deadlock Handling
In multi-robot social navigation, how an agent deals with deadlocks is very important. Deadlock handling can be divided into two main types: Prevention and Resolution. Generally, there are two main approaches: Proactive (Deadlock Prevention) and Reactive (Deadlock Resolution), each with its own style and advantages. The Proactive Strategy, on the other hand, involves planning ahead. Agents using this approach constantly analyze their environment to predict and prevent potential deadlocks before they happen. This is like a chess player who thinks several moves ahead. An example of this approach is IMPC-DR [6]. Proactive agents aim to avoid problems, ensuring smoother movement. However, this requires a good understanding and prediction of the environment, as wrong predictions can lead to unnecessary detours or inefficient paths. The Reactive Strategy focuses on immediate response. Agents using this method stay alert for potential deadlocks as they move. They don’t act until a deadlock is about to happen. Once they detect it, they quickly respond to resolve or avoid the deadlock. This strategy deals with problems as they come up. Examples of this approach are methods that include the right-hand rule [2], Auctions [5], and ORCA-MAPF [8]. The Reactive Strategy is flexible and can adapt to changes, but it might struggle if many problems occur at once. Fig. 9 and 10 categorize the MRNs for Deadlock Resolution and Prevention, respectively.
@@ III-C Properties of MRN in SMGs
III-C1 Safety
Provable safety can be achieved by single-integrator systems e.g. ORCA framework from Van Den Berg et al. [7] and its non-holonomic variant [55], which are effective for fast and exact multi-robot navigation. ORCA conservatively imposes collision avoidance constraints on the motion of a robot in terms of half-planes in the space of velocities. The optimal collision-free velocity can then be quickly found by solving a linear program. Proving safety is harder for systems with double-integrator dynamics, therefore safety in these systems depends on the planning frequency of the system. For example, the NH-TTC algorithm [148] uses gradient descent to minimize a cost function comprising a goal reaching term and a time-to-collision term, which rises to infinity as the agent approaches immediate collision. NH-TTC guarantees safety in the limit as the planning frequency approaches infinity. Other optimization-based approaches use Model Predictive Control(MPC); in such approaches, safety depends not only on the planning frequency but also on the length of the planning horizon. Finally, Control Barrier Functions (CBFs) [2] can be used to design controllers that can guarantee safety via the notion of forward invariance of a set i.e. if an agent starts out in a safe set at the initial time step, then it remains safe for all future time steps, that is, it will never leave the safe set.
III-C2 Liveness
Liveness is a property that means robots are continuously making progress towards their goals, and only stop once they have reached their goals. In general MRN, liveness is typically considered a “soft constraint” as opposed to safety, which is often a hard constraint. However, MRN in SMGs often result in deadlocks, so they become critical. Methods for ensuring liveness are characterized along the taxonomy can either be deadlock-preventing [22, 149, 150] or deadlock resolving [2, 138, 107, 143] (previously discussed in Sec. III-B3). The difference lies in when they solve a deadlock. Deadlock prevention methods are generally smoother and less invasive as they do not wait for a deadlock to happen in the first place. Next, certain liveness methods rely on agents communicating between themselves to resolve conflicts [97, 151, 98]. Finally, some liveness methods use global planning methods to plan non-colliding paths [8].
III-C3 Fairness

### v3

@@ Abstract
The “Last Mile Challenge” has long been considered an important, yet unsolved, challenge for autonomous vehicles, public service robots, and delivery robots. A central issue in this challenge is the ability of robots to navigate constrained and cluttered environments that have high agency (e.g., doorways, hallways, corridor intersections), often while competing for space with other robots and humans. We refer to these environments as “Social Mini-Games” (SMGs). Traditional navigation approaches designed for MRN do not perform well in SMGs, which has led to focused research on dedicated SMG solvers. However, publications on SMG navigation research make different assumptions (on centralized versus decentralized, observability, communication, cooperation, etc.), and have different objective functions (safety versus liveness). These assumptions and objectives are sometimes implicitly assumed or described informally. This makes it difficult to establish appropriate baselines for comparison in research papers, as well as making it difficult for practitioners to find the papers relevant to their concrete application. Such ad-hoc representation of the field also presents a barrier to new researchers wanting to start research in this area. SMG navigation research requires its own taxonomy, definitions, and evaluation protocols to guide effective research moving forward. This survey is the first to catalog SMG solvers using a well-defined and unified taxonomy and to classify existing methods accordingly. It also discusses the essential properties of SMG solvers, defines what SMGs are and how they appear in practice, outlines how to evaluate SMG solvers, and highlights the differences between SMG solvers and general navigation systems. The survey concludes with an overview of future directions and open challenges in the field. Our project is open-sourced at https://socialminigames.github.io/.
keywords:
Multi-Robot Navigation, Deadlocks, Social Navigation
1 Introduction
@@
(b) General multi-robot navigation (taken with permission from [15]).
@@
In recent years, the research field of MRN has started to coincide with the momentum of robot deployments in shared human spaces, largely driven by advances in data, compute, and foundation models. As a consequence, research has flourished in MRN, and more specifically, MRN in constrained environments that arise in shared human spaces. In Figure 1, we highlight two types of MRN in constrained scenarios. On the left, we depict Social Mini-Games (SMGs) [14, 16] on the right we depict “general MRN”. A key distinguishing factor between these scenarios is the notion of agency–that is, the capacity of an agent to take independent actions and influence the environment around them. SMGs exhibit high agency where small actions can significantly change the interaction dynamics, coordination requirements, and potential for deadlocks. General MRN, on the other hand, exhibit low to medium agency. High-agency SMGs have several unique challenges. First, without explicit scheduling, robots result in deadlocks, collisions, or non-smooth trajectories [17, 18]. Second, humans can avoid collisions and deadlocks without having to deviate too much from their nominal speed or trajectory. For instance, when two individuals go through a doorway together, one person modulates their velocity by just enough to enable the other person to pass through first, while still adhering closely to their preferred speed. This type of behavior presents a significant challenge for robots, especially in high-agency SMGs. Classical navigation approaches typically fail to cover all the aspects needed for SMGs [14]. So, new techniques, ideas, and systems are being investigated, often with roots deep within the larger parent field of MRN. This fact presents the crux of the issue: researchers studying SMG solvers from a MAPF perspective will build off of MAPF research, MARL-based SMG solvers will form their roots from the MARL community, and so on. Other communities build upon game theory and optimization. Each field has its own set of assumptions, limitations, objectives, and techniques. As a result, MRN in SMG research has fractured into these sub-communities without a clear, unified formulation. This categorization into sub-communities presents an entry barrier to new researchers in this field by making it difficult to navigate through and understand existing literature and to establish appropriate baselines for comparison. In addition, this makes it difficult for practitioners to find papers and solutions relevant to their concrete application. This article aims to address this growing challenge by introducing a unified taxonomy to describe MRN in SMG problems, and by establishing common assumptions, properties, and measures for evaluating MRN in SMG algorithms. The unified taxonomy we present in this paper is our attempt to classify the currently studied SMG solver variants. We hope this terminology will serve as a common ground for future researchers, and will be used by them to describe their contributions succinctly and accurately.
1.1 Differentiation from Existing Surveys
@@
Our survey focuses on multi-robot navigation in social mini-games, and thus belongs to the broader field of social robot navigation. Social robot navigation, however, is vast and we will not attempt to summarize it; instead, we refer to many recent surveys on social navigation [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Among these, [24] focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations on how to properly evaluate social navigation research. Another recent survey by Mavrogiannis et al. [25] focuses on the core challenges of social navigation with respect to different algorithms, human behavior models, and evaluation, but they focus on the general navigation case. Wang et al. [28] proposes new metrics evaluating the principles defined in [19], such as comfort, naturalness, and sociability. And finally, datasets [29, 30] and benchmarks [31, 32] have enabled researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. However, all of the surveys discussed above on various aspects of social robot navigation focus on robots in sparse or dense crowds, but not social mini-games. We showcase the difference in Figure 1. Perhaps the closest survey that mirrors the same objective as with this survey (that is, attempting to unify a growing, but fractured, research area) is the seminal survey on MAPF (“Multi-Agent Pathfinding: Definitions, Variants, and Benchmarks”) by Stern et al. [33]. Here, the authors present a definition for MAPF, describe the different variants of it, and introduce two benchmarks for proper evaluation. This paper was influential in anchoring and streamlining the future research in MAPF. We envision a similar success via our survey. The overall structure of this survey is given in Fig. 2. We first define the SMG problem in Sec. 2, highlight its various scenarios where it appears, and how to evaluate navigation approaches in SMGs. Then in Sec. 3, we introduce the taxonomy, properties, and algorithmic categories for research on SMG solvers. Then in Sec. 4, we benchmark and compare some recent state of the art SMG solvers on the scenarios introduced earlier. Finally in Sec. 5, we conclude by discussing general trends, open problems, and future directions.
@@
2 Social Mini-Games
2.1 SMGs versus General MRN
@@
We next give a mathematical definition of SMGs. It says that if we assume each agent has one or more individual optimal trajectory (for example, in the doorway scenarios, the optimal trajectory for both the green and the orange agent would be a straight line path towards the red cross.), then if all the individual optimal trajectories intersect for all agents within some time interval, then that scenario is an SMG.
@@
SMGs are embedded within MRN: most trajectories are decoupled, but at specific times (bottlenecks, crossings, occlusions) the interaction becomes a game with coupled best responses. MAPF also resolves conflicts but (classically) in discrete, fully observed, centralized settings that compute joint plans offline; SMGs arise online, continuous-time, often decentralized, with heterogeneous objectives and social norms. Furthermore, it is typical for SMGs to include around 2-4 agents in the most commonly occurring SMGs. Of course, an environment may consist of multiple SMGs occurring simultaneously. Our scenarios (Sec. 2.2) are chosen because they activate specific SMG characteristics, and our metrics (Sec. 2.3) include SMG-aware outcomes that are immaterial in decoupled MRN segments but decisive inside SMGs.
@@ 2.3 Evaluation Metrics
3 Algorithms, Taxonomy and Properties for SMGs
We analyze algorithms and system design choices conditioned on an active SMG, i.e., when at least one SMG property from Sec. 2.2 is present (bottleneck contention, crossing flows, persistent mutual occupancy, occlusion, or conflicting objectives). Let $\mathcal{K}(t)\!\subseteq\!\{1,\dots,N\}$ denote the active coupling set, that is, the minimal subset of agents whose best responses are mutually coupled during $[a,b]$ (Sec. 2.2). Coordination, communication, observability, invasiveness, cooperation, and deadlock handling are interpreted on $\mathcal{K}(t)$, not on the full set of agents. Throughout, we tie design choices to game-aware outcomes introduced in Sec. 2.3: Social Welfare Gap (SWG), Turn-Taking Fairness (TTF), and Invasiveness Score (IS), reported alongside standard MRN metrics. This section thus specifies how MRN methods become SMG solvers once a trigger activates and a coupling set emerges.
3.1 Categorization of SMG Solvers by Paradigm
@@
3.1.1 Multi-Agent Reinforcement Learning (MARL)
@@
3.1.2 Multi-Agent Path Finding (MAPF)
@@ 3.1.4 Others
3.2 Taxonomy of SMG Design Choices
@@ 3.2.1 Coordination
The last type of coordination is Decentralized which allows each agent to make its own decisions independently. Instead of relying on a central authority, each agent acts based on its own observations, rewards, and goals. This approach promotes independence and flexibility, enabling agents to adapt to their surroundings and other agents dynamically [71, 72]. Although it may seem less organized, with the right algorithms and planners, decentralized coordination can generate effective and efficient controls to develop a safe, collision-free, and deadlock-free trajectory [73, 36, 74, 75, 76]. Examples of this approach include IMPC-DR [77] and ORCA-MAPF [46], where agents independently navigate in their environment while avoiding collisions and deadlocks. Figures 4 and 5 list the key papers for the Centralized and Decentralized communication categories. Note that the color legend signifies the categorization based on Algorithmic/Methodological paradigm discussed in Sec. 3.1, where we use different colors for different approaches- multi-robot Reinforcement Learning (cyan), Optimization (violet), multi-robot Path Finding (green), and Others (Orange).
@@
[…]
Figure 6: Distributed Methods.     Multi-Agent RL    Optimization    MAPF    Others (e.g., Heuristic).
3.2.2 Communication
@@
3.2.3 Deadlock Handling
In multi-robot social navigation, how an agent deals with deadlocks is very important. Deadlock often refers to situations where robot states converge to an undesirable equilibrium that causes the robots to stall [103, 128, 129]. Deadlock handling can be divided into two main types: Prevention and Resolution. Generally, there are two main approaches: Proactive (Deadlock Prevention) and Reactive (Deadlock Resolution), each with its own style and advantages. The Proactive Strategy involves planning ahead. Agents using this approach constantly analyze their environment to predict and prevent potential deadlocks before they happen. This is like a chess player who thinks several moves ahead. Examples of this approach include IMPC-DR [77] and MERRY-GO-Round [130] that combine deadlock resolution strategies with trajectory planning and navigation to ensure smoother movements, thus avoiding potential deadlocks. However, this requires a good understanding and prediction of the environment, as wrong predictions can lead to unnecessary detours or inefficient paths. The Reactive Strategy, on the other hand , focuses on immediate response. Agents using this method stay alert for potential deadlocks as they move. They don’t act or react until a deadlock is about to happen. Once they detect it, the robots quickly respond to resolve or avoid the deadlock by taking temporarily adjusted trajectories. This strategy deals with problems as they come up. Examples of this approach are methods that include the right-hand rule [64], auctions [131], ORCA-MAPF [46], and adaptive rotational strategies [128, 129]. The Reactive Strategy is flexible and can adapt to changes, but it might struggle if many problems occur at once. Fig. 9 and 10 categorize the MRNs for Deadlock Resolution and Prevention, respectively.
[…]
@@ 3.3 Properties of MRN in SMGs
3.3.1 Safety
Provable safety can be achieved by single-integrator systems e.g. ORCA framework from Van Den Berg et al. [134] and its non-holonomic variant [48], which are effective for fast and exact multi-robot navigation. ORCA conservatively imposes collision avoidance constraints on the motion of a robot in terms of half-planes in the space of velocities. The optimal collision-free velocity can then be quickly found by solving a linear program. Proving safety is harder for systems with double-integrator dynamics, therefore safety in these systems depends on the planning frequency of the system. For example, the NH-TTC algorithm [153] uses gradient descent to minimize a cost function comprising a goal reaching term and a time-to-collision term, which rises to infinity as the agent approaches immediate collision. NH-TTC guarantees safety in the limit as the planning frequency approaches infinity. Other optimization-based approaches use Model Predictive Control(MPC); in such approaches, safety depends not only on the planning frequency but also on the length of the planning horizon. Finally, Control Barrier Functions (CBFs) [64] can be used to design controllers that can guarantee safety via the notion of forward invariance of a set i.e. if an agent starts out in a safe set at the initial time step, then it remains safe for all future time steps, that is, it will never leave the safe set. CBFs can also be extended to a probabilistic setting [154] to account for realistic uncertainties in system dynamics and observations.
3.3.2 Liveness
Liveness is a property that means robots are continuously making progress towards their goals, and only stop once they have reached their goals. In general MRN, liveness is typically considered a “soft constraint” as opposed to safety, which is often a hard constraint. However, MRN in SMGs often result in deadlocks, so they become critical. Methods for ensuring liveness are characterized along the taxonomy can either be deadlock-preventing [14, 155, 156] or deadlock resolving [64, 141, 103, 146] (previously discussed in Sec. 3.2.3). The difference lies in when they solve a deadlock. Deadlock prevention methods are generally smoother and less invasive as they do not wait for a deadlock to happen in the first place. Next, certain liveness methods rely on agents communicating between themselves to resolve conflicts [93, 157, 94]. Finally, some liveness methods use global planning methods to plan non-colliding paths [46].
3.3.3 Fairness